OpenStudy (anonymous):

Differentiation of the equation y = x^2 + z^2 with respect to x , thanks .

OpenStudy (anonymous):

2x

OpenStudy (aravindg):

y=2x+2z dz/dx

OpenStudy (anonymous):

@AravindG : Since he's calculating dy/dx, wouln't z just be a constant, therefore the derivative of it would be 0?

OpenStudy (aravindg):

Well it was not mentioned that z is a constant ..If it is the answer will simply be 2x

OpenStudy (anonymous):

@AravindG : Since he's calculating dy/dx, wouldn't it be assumed that Z is just a constant, since he's calculating partial derivative?

OpenStudy (kenljw):

y = x^2 + z^2 dy/dx = 2x or 2x +2xdz/dx

OpenStudy (kenljw):

2zdz/dx

OpenStudy (anonymous):

@FutureMathProfessor : Your concern should be a question , however I have started calculus after a long time break ; can you post a question with relevant example.

OpenStudy (anonymous):

@AravindG : Can you please give a proper explanation of PDE and ODE with regard to when to consider a variable and not a constant . Thanks.

OpenStudy (espex):

This question has been answered, what more do you need?

OpenStudy (anonymous):

@eSpeX -- Additional help , when to consider z a constant and not a variable .

OpenStudy (espex):

If you are asked to take a partial derivative of an equation containing more than one variable, you consider the other variables constants during the derivation of the other variables. If you are simply asked to take the derivative of an equation, it is proper to take the derivative of terms with other variables, with respect to x, and append that term with d<>/dx.

OpenStudy (espex):

Which calc are you in, you will be learning about partials towards the end of II and into III.

OpenStudy (anonymous):

An example ..

OpenStudy (anonymous):

@eSpeX : The last one basically .

OpenStudy (espex):

So you are in Calc III? I would expect that you will be covering this soon then.

OpenStudy (anonymous):

Well to be fair , I was given to solve this equation dy/dx = sqrt( y- x) , and it arised from that .

OpenStudy (espex):

An example using your equation. If you are asked to take the derivative of x^2 + z^2, you would write \[\frac{ dy }{ dx }=x^2 + z^2 \rightarrow 2x+2z\frac{dz}{dx}\] where as if you were asked to take the partial dz/dx and dy/dx of the same function, you would get \[\frac{dz}{dx}=x^2 + z^2\rightarrow 2z\]and\[\frac{dy}{dx}=x^2 + z^2\rightarrow 2x\]

OpenStudy (espex):

The 'd' would be different, I did not see a partial diff symbol in the equation library.

OpenStudy (anonymous):

I know about the 'd'

OpenStudy (anonymous):

Now last question , a little bit theoretical if you are interested , I read about it a long time back ; what is the difference between these two 'd's.

OpenStudy (espex):

As far as the origination I am uncertain.

OpenStudy (anonymous):

No its terms of application , the partial differentiation and the normal differentiation , just to brush up my memory .

OpenStudy (espex):

In terms of application they denote how you handle the function with multiple variables.

OpenStudy (kenljw):

y = x^2 + z^2 z = sqrt(y -x^2) why not the partial derivation dz/dx =-2x/sqrt(y - x^2)

OpenStudy (anonymous):

@KenLJW : We are not on the same page , may be .

OpenStudy (kenljw):

slight error dz/dx =-x/sqrt(y - x^2) was responding to previous answer for partial derivative